OpenStudy (anonymous):

find the perimeter of the half circle. Diameter is 40. show me the steps on how to get this.

OpenStudy (anonymous):

\[P_{o}=2 \pi r ~~~~or~~~~P_{o}=\pi d\]\[P_{o}=40 \pi\]\[P_{o}\approx 125.6\]\[half~of~total~perimeter={125.6\over2}\]\[\huge =62.8\]

OpenStudy (anonymous):

next you need to add the diameter

OpenStudy (anonymous):

i did that i got 102.83ft

OpenStudy (anonymous):

\[\huge 62.8+40=102.8\]

OpenStudy (anonymous):

Great! :D

OpenStudy (anonymous):

but my teacher posted a different answer for it, she must have been wrong

OpenStudy (anonymous):

thanks. i will question her about it tomorrow :)

OpenStudy (anonymous):

maybe you have to round to a certain place :)

OpenStudy (anonymous):

cool glad i could help :)

OpenStudy (anonymous):

no she gotv165.66ft

OpenStudy (anonymous):

oh...well then im not sure where she got that....be sure to question her to clear this up :)